1. Write an equation of the line containing the given point and perpendicular to the given line. Express your answer in the form y = mx + b
(4,5); 3x + y = 4
The equation of the line is y = What?
2. Write an equation of the line containing the given point and perpendicular to the given line.
(3, - 3); 8x +7y =
The equation of the line is y = What?

Answer by stanbon(75887)

): 1. Write an equation of the line containing the given point and perpendicular to the given line. Express your answer in the form y = mx + b
(4,5); 3x + y = 4
The equation of the line is y = What?
The given line has slope = -3.
So the equation you want must have slope = 1/3 and must pass thru (4,5)
y= mx+b
5 = (1/3)*4 + b
5 = 4/3 + b
15/3 - 4/3 = b
b = 11/3
EQUATION:
y = (1/3)x +(11/3)
------------------------------

2. Write an equation of the line containing the given point and perpendicular to the given line.
(3, - 3); 8x +7y = 0
The give line has slope = -8/7.
So the equation you want must have slope = 7/8 and pass thru (3,-3)
y = mx+b
-3 = (7/8)*3 + b
-3 = (21/8) + b
-24/8 = 21/8 + b
b = -45/8
-----------------
EQUATION:
y = (7/8)x - (45/8)
=========================
Cheers,
Stan H.